What mass of potassium nitrate is needed to generate 157.0 L of gas, composed of 105.0 L of N2 and 52.0 L of O2 at 0.920 atm and 293 K, using these two reactions?
1
Expert's answer
2020-10-23T09:15:34-0400

Nitrogen can be produced from potassium nitrate as indicated by the balanced equation:

10Na + 2KNO3 → K2O + 5Na2O + N2

PV = nRT

V(N2) = 105 L

P = 0.920 atm

T = 293 K

R = 0.08206 L×atm/mol×K

"n = \frac{PV}{RT}"

n(N2) "= \frac{0.920 \times 105}{0.08206 \times 293} = 4.02 \;mol"

According to the equation:

n(KNO3) = 2n(N2) "= 2 \times 4.02 = 8.04 \;mol"

M(KNO3) = 101.1 g/mol

"m = n\times M"

m(KNO3) "= 8.04 \times 101.1 = 812.84 \;g"

Oxygen gas can be generated by heating potassium chlorate as indicated by the balanced equation:

2KClO3 → 2KCl + 3O2

V(O2) = 52.0 L

n(O2) "= \frac{0.920 \times 52}{0.08206 \times 293} = 1.99 \;mol"

According to the equation:

n(KClO3) = 2n(O2) "= 2 \times 1.99 = 3.98 \;mol"

M(KClO3) = 122.55 g/mol

m(KClO3) "= 3.98 \times 122.55 = 487.75 \;g"